# SafetyAnalyst: Interpretable, transparent, and steerable LLM safety moderation

### v1

@@ 3 Evaluating SafetyReporter against existing LLM safety moderation systems on prompt classification
To evaluate the effectiveness of SafetyAnalyst on identifying potentially harmful prompts, we tested SafetyReporter (aligned to WildJailbreak prompt labels with weights illustrated in Figure 3) on a comprehensive set of public benchmarks featuring potentially unsafe user queries and instructions against existing LLM safety moderation systems. Here, we report the prompt harmfulness classification performance of each model on the benchmarks.
3.1 Evaluation setup
Benchmarks.
We tested SafetyReporter and relevant baselines on 6 publicly available prompt safety benchmarks, including SimpleSafetyTests (100 prompts; Vidgen et al. 2023), HarmBenchPrompt standard test set (159 prompts; Mazeika et al. 2024), WildGuardTest (960 vanilla and 796 adversarial prompts; Han et al. 2024), AIR-Bench-2024 (5,694 prompts; Zeng et al. 2024c), and SORRY-Bench (9,450 prompts; Xie et al. 2024). These benchmarks represent a diverse and comprehensive selection of unsafe prompts, including manually crafted prompts on highly sensitive and harmful topics (SimpleSafetyTests), standard behavior that may elicit harmful LLM responses (HarmBench), adversarial prompts (WildGuardTest), benign prompts (WildGuardTest), prompts that may challenge government regulations and company policies (AIR-Bench-2024), and unsafe prompts that cover granular risk topics and linguistic characteristics (SORRY-Bench). Since our system focuses on identifying prompts that would be unsafe to respond to, rather than the harmfulness in the prompt content per se, we did not include benchmarks in which prompts were labeled for the latter, such as the OpenAI moderation dataset (Markov et al., 2023), ToxicChat (Lin et al., 2023), and AegisSafetyTest (Ghosh et al., 2024).
Baselines.
We compare SafetyReporter to 9 existing LLM safety moderation systems: OpenAI moderation endpoint (Markov et al., 2023), LlamaGuard, LlamaGuard-2, LlamaGuard-3 (Inan et al., 2023), Aegis-Guard-Defensive, Aegis-Guard-Permissive (Ghosh et al., 2024), ShieldGemma-2B, ShieldGemma-9B, ShieldGemma-27B (Zeng et al., 2024a), and WildGuard (Han et al., 2024). Additionally, we report zero-shot GPT-4 performance (Achiam et al., 2023).
We referenced Han et al. (2024)’s evaluation results where applicable and additionally tested models and benchmarks that they did not feature with temperature set to 0. We were unable to fairly evaluate Llama-Guard, Aegis-Guard-Defensive, and Aegis-Guard-Permisive (both Aegis-Guards are tuned Llama-Guard models) on SORRY-Bench, since the lengths of 457 prompts in SORRY-Bench exceeded the Llama-2 context window limit of 4,096 tokens (Touvron et al., 2023). For each model, we computed an average F1 score across benchmarks weighted by the number of prompts in each benchmark dataset.
3.2 Evaluation results
SafetyReporter outperforms existing LLM safety moderation systems on prompt harmfulness classification.
Table 3 shows our evaluation results, measured by the F1 score (denoted in percentage). SafetyReporter achieved competitive performance on all benchmarks compared to existing LLM safety moderation systems, with the highest overall F1 score of 75.4, exceeding the second highest score of 71.7 by WildGuard. Notably, SafetyReporter’s performance was zero-shot, since it was not trained on or aligned to any training datasets of the benchmarks, whereas WildGuard was trained on the WildGuardTrain set. Nonetheless, GPT-4’s classification performance was better than all the LLM moderation models with an F1 score of 81.6.
@@ 3.2 Evaluation results
Model SimpSTests HarmBench WildGuardTest AIR-Bench SORRY-Bench Average
Vani. Adv.
@@
SafetyReporter 95.2 94.4 88.3 73.7 83.0 69.1 75.4
@@ 4 Related work
Existing LLM content moderation systems.
@@
Existing LLM content moderation systems include WildGuard (Han et al., 2024), ShieldGemma (Zeng et al., 2024a), AegisGuard (Ghosh et al., 2024), LlamaGuard (Inan et al., 2023), and the OpenAI moderation endpoint (Markov et al., 2023). These systems are LM-based classifiers that can categorize content risk, including user prompts. Although they can achieve high classification accuracy on prompt safety benchmarks (e.g., classifying a prompt as harmful or benign), their internal decision mechanisms are challenging to interpret, which limits their reliability and generalizability. Furthermore, due to the lack of modularity in their architectures, they cannot be easily steered to reflect different safety perspectives.
LLM content risk.
@@
The AI safety literature has relied on risk taxonomies to categorize unsafe content. Recent work has built on standard risk categories (Weidinger et al., 2022) to include more fine-grained categories (Wang et al., 2023; Tedeschi et al., 2024; Xie et al., 2024; Brahman et al., 2024), achieve comprehensive coverage (Vidgen et al., 2024), and incorporate government regulations and company policies (Zeng et al., 2024b). Our system relies on the taxonomy developed by Zeng et al. (2024b), selected for its comprehensive and fine-grained nature. Overall, these taxonomies describe the unsafe nature of a prompt or unsafe actions that might result from a prompt being answered. To our knowledge, no prior work exists that proposes formal taxonomies for the downstream effects of unsafe prompts (as opposed to actions; see Appendix A for our taxonomies of harmful and beneficial effects).
Symbolic knowledge distillation.
We distilled a pair of small, expert LMs (SafetyReporter) to create structured harm-benefit trees, the core of our interpretable framework. The symbolic knowledge distillation strategy leverages diffuse knowledge gained by large, generalist (and often proprietary) models to create a more compact expert student model that excels at one particular task (Xu et al., 2024; West et al., 2021; Tang et al., 2019). This strategy is useful (among other reasons) to generate rich, structured data that is too costly or labor-intensive for humans to do by hand (West et al., 2021). Indeed, prior work shows that symbolic knowledge distillation from machine teachers can exceed the quality of human-authored symbolic knowledge (West et al., 2021; Jung et al., 2023). Compared to the teacher models, our SafetyReporter uses less time, memory, compute, and cost while achieving comparable performance, and it will be openly released for public use in LLM moderation contexts.
Pluralistic alignment for LLM safety.
@@
Although current LLM safety moderation systems are yet to be pluralistically aligned, recent interest in value pluralism Sorensen et al. (2024a) has given rise to rapid developments of pluralistic alignment approaches for LLMs. Lera-Leri et al. (2022) formalized an aggregation method for value systems inspired by the social choice literature. Feng et al. (2024) outlined a more general framework based on multi-LLM collaboration, in which an LLM can be aligned to specialized community LMs for different pluralism objectives. Other methods have been proposed for learning distributions of human preferences rather than the majority (Siththaranjan et al., 2023; Chen et al., 2024).
5 Discussion
In this paper, we introduce SafetyAnalyst, an interpretable, transparent, and steerable framework for LLM content safety moderation. In addition to the conceptual framework, we provide an implementation of the system with various resources: a large-scale dataset of rich safety features (in the form of a structured “harm-benefit tree”) generated by SOTA LLMs on 19k prompts, the open-source SafetyReporter for harm-benefit feature generation, the first taxonomies of harmful and beneficial effects for AI safety, and a feature aggregation algorithm that can be steered to align with a given safety content label distribution or with top-down safety standards.
Our application of SafetyAnalyst and SafetyReporter to a comprehensive set of prompt safety benchmarks shows SOTA performance compared to existing LLM safety moderation systems. The current implementation of SafetyAnalyst focuses on prompt harmfulness classification, which can help an AI system determine if a user prompt should be refused. However, this framework can be extended to solve other content safety tasks, such as LLM response moderation, free-text moderation (Zhang et al., 2023), and general text moderation.
Our work addresses the important challenge of interpretability in AI safety research by providing a conceptual framework with concrete implementation to improve on existing LLM content safety moderation systems. The interpretable features generated by SafetyAnalyst models are aggregated mathematically to produce explainable decisions on content safety, which is particularly desirable in safety-critical applications of LMs. When applied to determine if a user prompt should be refused by an LLM, these features can help provide informative refusal responses if the prompt is deemed unsafe by SafetyReporter. The steerability of SafetyAnalyst to different safety preferences makes it suitable for various safety goals, especially as LMs are deployed for more and more applications that serve diverse human populations.
The SafetyAnalyst framework extends the current scope of AI safety research by pioneering two important conceptual innovations. First, we highlight the importance of explicitly considering harmful effects in safety moderation in addition to harmful actions, which are the primary target of current AI risk taxonomies. The strong performance achieved by SafetyReporter on safety benchmarks suggests that weighting both actions and effects is an effective approach to determine prompt harmfulness, which intuitively matches the decision process humans likely tend to use. Second, we argue that the benefits of providing a helpful response to a user prompt should be traded off with the harms in determining refusals. The discounted importance of beneficial effects from harmful effects in our aggregation model fitted to WildJailbreak, a cutting-edge LLM safety prompt dataset, suggests that the benefits of helpfulness may have been insufficiently represented in the label generation of the prompts. Future prompt harmfulness benchmarks and safety systems should account for effects and benefits in addition to only harmful actions to achieve more robust safety properties.
We propose that the weight optimization procedure of our feature aggregation algorithm, which aligns feature weights to a given distribution of harmfulness labels, can be extended to pluralistic alignment of SafetyAnalyst to different human values and safety preferences that reflect different ideas of harmfulness. Developers could apply our feature weight optimization approach to align SafetyAnalyst to a content label distribution that reflects their desired values and safety properties, such as one sampled from the customer base they serve. Although we weighted action categories at level-2 of the AIR 2024 taxonomy (e.g., Hate/Toxicity), our aggregation algorithm can be expanded to optimize weights on more fine-grained, level-3 categories (e.g., Harassment, Hate Speech, Perpetuating Harmful Beliefs, and Offensive Language) due to the taxonomy’s extensive coverage of AI risk. Furthermore, the aggregation algorithm can be extended to additionally parameterize the relative importance of effect categories that may vary between human populations and LM applications with different safety goals.
Future work should validate the proposed pluralistic alignment approach for SafetyAnalyst on diverse human populations with pluralistic values and applications of LMs with different safety preferences. Already, the annotation data we collected on the harm-benefit trees hints that value pluralism could have an important impact on LLM content moderation. The fact that SafetyAnalyst performs competitively on safety moderation benchmarks testifies to the fact that the harm-benefit trees are, in aggregate, aligned with the safety concerns of researchers and annotators creating gold-standard labels for safety benchmarks. However, the results in Table 5 reveal a more complex picture. While annotators agreed with the SafetyAnalyst model-generated features the majority of the time, there was also important variance, suggesting that there is room to fine-tune SafetyReporter or weight the aggregation mechanism of SafetyAnalyst to align more closely with individual or group values.
Limitations.
Generating the extensive harm-benefit trees, which are crucial to the interpretability of SafetyAnalyst, leads to longer inference time compared to existing, less interpretable LLM moderation systems. Although our specialized SafetyReporter substantially reduces the cost of feature generation than using an off-the-shelf LLM, we make the conscious trade-off between interpretability and efficiency to make LLM content safety decisions more reliable and transparent. While our system draws on the principles of cost-benefit-analysis commonly used to justify the adoption of governmental policies, following Arrow et al. (1996) we emphasize that simply summing harmful and beneficial effects will not be ultimately sufficient for safe decision-making. Future work should explore issues related to the incommensurability of values, the effectiveness with which SafetyAnalyst captures non-quantifiable harms and benefits, and the importance of weighting actions themselves, beyond just the effects they produce.
Conclusion.
@@
We introduce SafetyAnalyst, a novel conceptual framework for interpretable, transparent, and steerable LLM content safety moderation. We operationalized the pipeline of harm-benefit tree data generation, symbolic knowledge distillation, and weighted feature aggregation to implement an LLM safety moderation system for prompt harmfulness classification. Our system achieved SOTA performance on a comprehensive set of prompt safety benchmarks, which promises strong potential in real-world LLM safety applications.
Appendix A Prompting and definitions for harm-benefit feature generation

### v3

@@ 3.1 Prompt Safety Classification
To evaluate the effectiveness of SafetyAnalyst on identifying potentially harmful prompts, we tested it (with the aggregation model aligned to WildJailbreak and weights illustrated in Figure 3) on a comprehensive set of public benchmarks featuring potentially unsafe user queries and instructions against existing LLM content safety moderation systems. Here, we report the prompt harmfulness classification performance of each model on every benchmark.
@@
Model SimpSTests (n=100) HarmBench (n=159) WildGuardTest AIR-Bench (n=5,694) SORRY-Bench (n=9,450) Average
Vani. (n=960) Adv. (n=796)
@@
SafetyAnalyst 88.8 96.1 90.9 79.6 88.9 75.4 81.2
Benchmarks
We tested SafetyAnalyst and relevant baselines on 6 publicly available prompt safety benchmarks, including SimpleSafetyTests (100 prompts; Vidgen et al. 2023), HarmBenchPrompt standard test set (159 prompts; Mazeika et al. 2024), WildGuardTest (960 vanilla and 796 adversarial prompts; Han et al. 2024), AIR-Bench-2024 (5,694 prompts; Zeng et al. 2024c), and SORRY-Bench (9,450 prompts; Xie et al. 2024). These benchmarks represent a diverse and comprehensive selection of unsafe prompts, including manually crafted prompts on highly sensitive and harmful topics (SimpleSafetyTests), standard behavior that may elicit harmful LLM responses (HarmBench), adversarial prompts (WildGuardTest), benign prompts (WildGuardTest), prompts that may challenge government regulations and company policies (AIR-Bench-2024), and unsafe prompts that cover granular risk topics and linguistic characteristics (SORRY-Bench). Since SafetyAnalyst focuses on identifying prompts that would be unsafe to respond to, rather than the harmfulness in the prompt content per se, we did not include benchmarks in which prompts were labeled for the latter, such as the OpenAI moderation dataset (Markov et al., 2023), ToxicChat (Lin et al., 2023), and AegisSafetyTest (Ghosh et al., 2024).
Baselines
We compare SafetyAnalyst to 9 existing LLM safety moderation systems: OpenAI moderation endpoint (Markov et al., 2023), LlamaGuard, LlamaGuard-2, LlamaGuard-3 (Inan et al., 2023), Aegis-Guard-Defensive, Aegis-Guard-Permissive (Ghosh et al., 2024), ShieldGemma-2B, ShieldGemma-9B, ShieldGemma-27B (Zeng et al., 2024a), and WildGuard (Han et al., 2024). Additionally, we report zero-shot GPT-4 performance (Achiam et al., 2023). Appendix C contains detailed descriptions of all baselines evaluated. We referenced Han et al. (2024)’s evaluation results where applicable and additionally tested baselines and benchmarks that they did not feature, with temperature fixed to 0 for deterministic and reproducible results. We were unable to fairly evaluate Llama-Guard, Aegis-Guard-Defensive, and Aegis-Guard-Permissive (both Aegis-Guards are tuned Llama-Guard models) on SORRY-Bench, since the lengths of 457 prompts in SORRY-Bench exceeded the Llama-2 context window limit of 4,096 tokens (Touvron et al., 2023). For each model, we computed an average F1 score across benchmarks weighted by the number of prompts in each benchmark dataset. Experiments using open-weight models were run on one NVIDIA H100 GPU with batched inference using vllm (Kwon et al., 2023).
Results
Table 2 shows our evaluation results, measured by the F1 score (denoted in percentage). SafetyAnalyst achieved competitive performance on all benchmarks compared to existing LLM safety moderation systems, with the highest overall F1 score of 81.2, exceeding the second highest score of 71.7 by WildGuard. Nonetheless, GPT-4’s classification performance was marginally better than SafetyAnalyst’s, with an F1 score of 81.6. In Appendix C.3, we show that GPT-4’s outstanding performance on SORRY-Bench was driven by its better capability to identify potentially unsafe prompts encoded or encrypted in Atbash and Caesar ciphers, and that SafetyAnalyst outperformed other baselines in identifying potentially unsafe prompts against Persuasion Techniques (Authority Endorsement, Evidence-based Persuasion, Expert Endorsement, Logical Appeal, and Misrepresentation). Although SafetyAnalyst achieved outstanding performance in our evaluation, there remains room for improvement. In Appendix F, we highlight two main types of failure cases and discuss potential approaches to address them.
3.2 Inference Cost
Due to the extensiveness of the harm-benefit trees generated by SafetyAnalyst for each prompt (Figure 1; Table 4), it requires more inference-time compute than other baselines that only produce safety labels. In our evaluations, SafetyAnalyst averaged 6.12 seconds per prompt, which is longer than 0.22 seconds per prompt for WildGuard, the second-best baseline, with the same hardware configuration (one H100 GPU). Therefore, SafetyAnalyst is best reserved for cases where interpretable, robust, and steerable safety moderation is valued over compute usage at inference time (Zaremba et al., ).
To minimize the inference cost of SafetyAnalyst, we distilled the harm-benefit tree generation capability from frontier teacher LMs into a pair of lightweight, specialized student LMs—substantially reducing both the financial and computational costs (Table 5) without compromising performance (Table 1). Additionally, the separation of harm-tree and benefit-tree generation into two student models allows them to execute in parallel, thereby reducing the inference time without parallelization (6.12 seconds per prompt) by half. Note that it is possible that SafetyAnalyst’s inference could be further accelerated by parallel computing, which we may not have fully optimized here.
Finally, if a faster instantiation of SafetyAnalyst were desired, a promising approach would be to selectively ablate different types of features in the harm-benefit trees to only preserve the most helpful ones. As a demonstration of this approach, we systematically ablated different dimensions of the harm-benefit trees and report SafetyAnalyst’s performance on WildGuardTest and WildJailbreak (Appendix C.5). Our results show that harms contributed more than benefits, and likelihood more than extent and immediacy in the aggregation algorithm aligned to WildJailbreak. However, since this observation may not hold true for all datasets and tasks (particularly for those where disagreements among annotators are likely), we report the full harm-benefit tree in the current work for generality.
3.3 Harm-Benefit Feature Quality
To evaluate the quality of generated harm-benefit features, we collected human annotation data from 126 Prolific workers on their agreement with the generated stakeholders, harmful/beneficial effects, and the likelihoods, extents, and immediacies of the effects. Annotators showed broad agreement on the plausibility of the harm-benefit features (see Table 7 in Appendix D for results and Figure 5 in Appendix D for interface design).
4 Additional Benefits of SafetyAnalyst
Interpretability and Transparency
Although SafetyAnalyst achieved outstanding performance on prompt safety classification, its most critical advantage is the interpretability of its decision-making process compared to black-box systems, including all the baselines in Table 2. This interpretability is two-fold: first, the features, on which the safety predictions are based solely, are explicitly generated by SafetyAnalyst and semi-structured (i.e., on carefully curated dimensions including stakeholder, harm, benefit, action, effect, extent, likelihood, and immediacy); second, these features are aggregated using a white-box model with transparent mechanisms and interpretable feature weights that quantify the importance of corresponding feature values (Figure 3). Even though LLMs (such as GPT-4) can generate explanations for their decisions, there still lacks interpretability in how the decisions are reached and reliable causal relationships between the explanation and safety prediction. Our strong evaluation results in Table 2 suggest that our simple but interpretable features and aggregation model contain sufficient information for making prompt safety predictions. Appendix E includes a detailed example of the full decision-making process of SafetyAnalyst, highlighting its interpretability and transparency.
Steerability
In addition, SafetyAnalyst’s aggregation model is defined by a set of transparent, interpretable weight parameters. The parameters aligned to WildJailbreak we report in Figure 3 measure the internal safety values of the annotators or LMs who provided the labels for the WildJailbreak dataset. However, one central strength of SafetyAnalyst is that the aggregation model allows different safety features to be up- or down-weighted for top-down adjustments, or aligned to a customized safety label distribution for bottom-up adjustments. We provide concrete explanations for how to operationalize top-down weight adjustments in the case study in Appendix E. Bottom-up adjustments of weights can be achieved by optimizing the aggregation model parameters to a safety label distribution produced by an individual or group; the resulting parameters would be aligned to the safety values and preferences inherently expressed in the annotated labels.
@@ 5 Related Work
AI Safety Moderation
@@
AI safety moderation refers to the process of ensuring that AI systems operate safely, ethically, and in alignment with human values (Han, 2024; Achara & Chhabra, 2025). Main approaches to AI safety include red-teaming (Lin et al., 2024), content moderation (Huang et al., 2024), privacy (Korobenko et al., 2024), alignment (Shen et al., 2024), security (Qi et al., 2024), and governance (Birkstedt et al., 2023). Our SafetyAnalyst framework addresses content moderation, with an emphasis on AI behavior, while the system presented in this paper specializes in moderating LLM prompts. Existing LLM content moderation systems include WildGuard (Han et al., 2024), ShieldGemma (Zeng et al., 2024a), AegisGuard (Ghosh et al., 2024), LlamaGuard (Inan et al., 2023), and the OpenAI moderation endpoint (Markov et al., 2023). These systems are LM-based classifiers that can categorize content risk, including user prompts. Except for minor variations, each of these systems is structured similarly: a general-purpose LLM is trained on a large dataset that links contents (e.g., prompts) to safety labels. The resulting content moderation systems can then classify some content as harmful or not based on the training they received (see Appendix C.1 for details). Although some systems built in this way can achieve high classification accuracy on prompt safety benchmarks (e.g., classifying a prompt as harmful or benign), their internal decision mechanisms are challenging to interpret (there is no straightforward way to determine why a prompt was classified as harmful by one of these systems), which limits their reliability and generalizability. Furthermore, due to the lack of modularity in their architectures, they cannot be easily steered to reflect different safety perspectives beyond expensive and time-consuming retraining or fine-tuning processes.
AI Safety Taxonomies
@@
Prior work has characterized AI safety based on the potential of risk and, therefore, relied on risk taxonomies to categorize unsafe content and behavior (Bai et al., 2022; Shen et al., 2023; Huang et al., 2024; Ji et al., 2024). Recent work has built on standard risk categories (Weidinger et al., 2022) to include more fine-grained categories (Wang et al., 2023; Tedeschi et al., 2024; Xie et al., 2024; Brahman et al., 2024), achieve comprehensive coverage (Vidgen et al., 2024), and incorporate government regulations and company policies (Zeng et al., 2024b). Our system relies on the taxonomy developed by Zeng et al. (2024b), selected for its comprehensive and fine-grained nature. In the context of our work, these taxonomies describe the unsafe nature of a prompt or unsafe actions that might result from a prompt being answered. To our knowledge, no prior work exists that proposes formal taxonomies for the downstream effects of unsafe prompts (as opposed to actions; see Appendix A for our taxonomies of harmful and beneficial effects).
Knowledge Distillation
Knowledge distillation aims to transfer the capabilities of large, complex “teacher” models to smaller, more efficient “student” models. Existing techniques transfer knowledge from output logits (Hinton et al., 2015), intermediate layer representations (Romero et al., 2014), attention representations (Zagoruyko & Komodakis, 2017), or symbolic knowledge (West et al., 2021). Due to the proprietary nature of some teacher models, we applied symbolic knowledge distillation to train the student model, which enabled us to transfer knowledge solely via model outputs and data without the need to access internal representations of teacher models.
Pluralistic Alignment
@@
Although current AI safety moderation systems are yet to be pluralistically aligned, recent interest in value pluralism (Sorensen et al., 2024a) has given rise to rapid developments of pluralistic alignment approaches for LLMs. Lera-Leri et al. (2022) formalized an aggregation method for value systems inspired by the social choice literature. Feng et al. (2024) outlined a more general framework based on multi-LLM collaboration, in which an LLM can be aligned to specialized community LMs for different pluralism objectives. Other methods have been proposed for learning distributions of human preferences rather than the majority (Siththaranjan et al., 2023; Chen et al., 2024). Additionally, some recent work has featured individualized human preference data, including the DICES dataset (Aroyo et al., 2024) and the PRISM alignment project (Kirk et al., 2024), paving the path to pluralistically or personally aligned AI systems.
6 Conclusion
@@
We introduce SafetyAnalyst, a novel conceptual framework based on LM-generated, semi-structured harm-benefit trees for interpretable, transparent, and steerable safety moderation for AI behavior. We operationalized the pipeline of harm-benefit tree data generation through chain-of-thought prompting, symbolic knowledge distillation, and weighted feature aggregation to implement a system for LLM prompt safety classification. Our system achieved SOTA performance on a comprehensive set of prompt safety benchmarks, promising strong potential in real-world LLM safety applications.
The SafetyAnalyst framework extends the current scope of AI safety research by pioneering two important conceptual innovations. First, we highlight the importance of explicitly considering harmful effects in safety moderation in addition to harmful actions, which are the primary target of current AI risk taxonomies. The strong performance achieved by SafetyAnalyst on safety benchmarks suggests that weighting both actions and effects is an effective approach to determine prompt harmfulness, which intuitively matches the decision process humans likely tend to use. Second, we argue that the benefits of AI behavior should be traded off with the harms. The discounted importance of beneficial effects from harmful effects in our aggregation model aligned to WildJailbreak, a cutting-edge LLM safety prompt dataset, suggests that the benefits of helpfulness may have been insufficiently represented in the ground-truth labeling of the prompts. Future prompt safety benchmarks and systems should account for effects and benefits in addition to only harmful actions to achieve more robust safety properties.
We propose that our aggregation model weight optimization procedure, which aligns the weights to a given label distribution, can be extended to pluralistic alignment of SafetyAnalyst to different safety preferences and community values that reflect different ideas of what constitutes harmfulness. Developers could apply our weight optimization approach to align an implementation of SafetyAnalyst to some label distribution that reflects their desired values and safety properties, such as one sampled from the customer base they serve.
Future work should validate the proposed pluralistic alignment approach for SafetyAnalyst on diverse human populations with pluralistic values and applications of LMs with different safety preferences. Already, our annotation data on the harm-benefit trees hints that value pluralism could have an important impact on LLM content moderation. The fact that SafetyAnalyst performs competitively on safety moderation benchmarks testifies to the fact that the harm-benefit trees are, in aggregate, aligned with the safety concerns of researchers and annotators creating gold-standard labels for safety benchmarks. However, the results in Table 7 reveal a more complex picture. While annotators agreed with the LM-generated features the majority of the time, there was also important variance, suggesting that there is room to fine-tune SafetyAnalyst or the aggregation model to align more closely with individual or group values.
Limitations
Generating the extensive harm-benefit trees, which are crucial to the interpretability of the SafetyAnalyst framework, leads to longer inference time compared to existing, less interpretable LLM moderation systems. Although our distilled student LMs substantially reduce the cost of feature generation (compared to using a frontier LLM), we make the conscious trade-off between interpretability and efficiency to make LLM content safety decisions more reliable and transparent. While our system draws on the principles of cost-benefit analysis commonly used to justify the adoption of governmental policies, following Arrow et al. (1996), we emphasize that simply summing harmful and beneficial effects will not be ultimately sufficient for safe decision-making. Future work should explore issues related to the incommensurability of values, the effectiveness with which SafetyAnalyst captures non-quantifiable harms and benefits, and the importance of weighting actions themselves, beyond just the effects they produce.
Impact Statement
This paper presents work whose goal is to advance the field of AI safety moderation by introducing a framework that emphasizes interpretability, transparency, and steerability. While the primary focus is on improving the reliability and trustworthiness of AI systems, we acknowledge the broader societal implications of our work. By enabling more transparent and customizable safety moderation, our framework has the potential to enhance the alignment of AI systems with diverse human values and safety standards. This could lead to more ethical AI deployments, particularly in sensitive applications where understanding and moderating AI behavior is critical. However, as with any advancement in AI, there are potential risks, such as the misuse of steerable systems to align with harmful or biased preferences. We encourage ongoing dialogue and ethical scrutiny to ensure that such systems are developed and deployed responsibly. In this context, we believe the societal consequences of our work are largely positive, but we remain vigilant about potential unintended outcomes.
Author Contributions
J.-J. L., V. P., M. K.-W., Y. C., and S. L. contributed to the conceptualization of the research. J.-J. L. performed experimentation, analysis, and manuscript writing under S. L.’s supervision. V. P., M. K.-W., L. J., N. D., A. G. E. C., J. S. B., M. S., and Y. C. provided valuable feedback and guidance. Y. C. and S. L. managed funding acquisition.
Acknowledgments
This work was supported by the Templeton World Charity Foundation Award Number TWCF-2023-32585, Cooperative AI Foundation, and Foresight Institute.
Appendix A Prompting Scheme and Taxonomies